OpenStudy (anonymous):

save me? Final exam coming up need help on number 24 in this packet

OpenStudy (anonymous):

number 24 need help please

OpenStudy (anonymous):

and apparently there is no use of a calculator

OpenStudy (hba):

Are you gonna fail ?

OpenStudy (anonymous):

lol no why you say that?

OpenStudy (anonymous):

Which letter are you stuck on?

OpenStudy (anonymous):

just in general... how should i go about it? i dont want to just go plugging in numbers from 0-24 and waste all that time if there is a simpler way of doing it

OpenStudy (anonymous):

Do you know how to find the max/min of a function?

OpenStudy (anonymous):

care to remind me? lol

OpenStudy (anonymous):

Take the derivative of the function and set it equal to zero

OpenStudy (anonymous):

could you explain that by showing me how you'd find the answer to a?

OpenStudy (anonymous):

Yeah first take the derivative of T(t) and let me know what you get

OpenStudy (anonymous):

we havent dont derivatives yet... this is precalc lol

OpenStudy (phi):

the sin function goes between -1 and 1 if you have \[ A\sin(\omega t + \phi) \] \(A\) is the amplitude, \(\omega\) is the angular frequency = \( 2 \pi f\) and ϕ is the phase, or in terms of period T \[ A\sin(\frac{2\pi}{T} t + \phi) \]

OpenStudy (phi):

The max value of your expression must occur when sin is at a max value of 1 Similarly, the min value occurs at sin() = -1

OpenStudy (anonymous):

so how would i set up this equation to make it so that it gives me the max and min? sorry or being blunt but my final is in an hour lol

OpenStudy (phi):

for part (d) plug in t= 12 and for (e) plug in t=18 for the period, match up \[ \frac{\pi}{12}= \frac{2 \pi}{T} \] and solve for T you get T= 24 hours

OpenStudy (phi):

so how would i set up this equation to make it so that it gives me the max and min? The equation is 13 sin(stuff) + 62 the only thing that changes is the value of sin(stuff). sin has a max of 1, so the max of the equation is 13*1+62= 75 the min of sin() is -1, so the min of the whole expression is 13*-1+62= 62-13= 49

OpenStudy (anonymous):

lol you're a beast, got it =) and how did you get d and e? sorry

OpenStudy (anonymous):

and c, when it tells us to plug in, how do we calculate that?

OpenStudy (phi):

for (d) replace t with 12 \[ 13\sin \left(\frac{\pi}{12} \cdot 12 - \frac{2\pi}{3} \right) + 62 \] They expect you to memorize sin, cos and tan of 0, 30, 45, 60 and 90 or in radians 0, pi/6, pi/4, pi/3, pi/2 and the sign according to which quadrant (for angles greater than 90)

OpenStudy (phi):

if you have a calculator, you can use it to refresh your memory...

OpenStudy (phi):

**(c) not (d). for (d) use t= 18 (6 hours past 12 noon)

OpenStudy (anonymous):

so is it 13 sin (90-270)+62?

OpenStudy (phi):

if want to use degrees multiply by 180/pi \[ 13\sin \left(\frac{\pi}{12} \cdot 12 - \frac{2\pi}{3} \right) + 62 \] pi/12*12 becomes pi or just 180 2pi/3 becomes 2 pi /3 *180/pi = 120 180-120= 60 degrees \[ 13\sin \left(60º \right) + 62 \]

OpenStudy (phi):

and sin(60) is sqrt(3)/2

OpenStudy (anonymous):

got it thanks a lot man you were great

OpenStudy (phi):

good luck